The lentgh of a rectangle is 2in. more than than twice its width . If the perimeter of the rectangle is 34 in.,find the dimensions of type rectangle.
Answer by nadinadan(32)

First let's say L=length and W=width of the rectangle.
The problem has to be changed into mathematical operations.
So it says that L=2W+2 length is 2plus double the width.
And the perimeter is 34. The perimeter of the rectangle is 2L+2W=34
so we have 2 equations:
2L+2W=34
L=2W+2 we will replace the second equation into the first one.
2(2W+2)+2W=34 so 4W+4+2W=34 so 6W=34-4 so 6W=30 so W=5
Now we can find what L is L=2W+2=2*5+2=12 so L=12
We can verify ourselves: P=2L+2W=24+10=34 YEAP, we are right.
Let me know if you need more help:
nadinadan@hotmail.com